c. To 238.00 milliliters of the buffer in part a, you add 13.38 mL of 2.500 M HClO!. What is the pH after that addition? Yes, you may use the H-H equation here. For your convenience, the buffer is 0.125 M formic acid (?! = 1.77 x 10!!) and 0.140 M sodium formate. (6 pts.)
Answer #1

pt - pka_ -log [ka =-log (1.8x109 Thcool pH pka + log (HCosti] = 3.75 +log [ 0.14] - 3.75 [0.125] = 3.75 +0.0492 = = 3. 7992